OpenStudy (anonymous):

Which pair has equally likely outcomes? List the letters of the two choices below which have equal probabilities of success, separated by a comma. A standard deck of cards has 12 face cards and four Aces (Aces are not face cards). A. rolling a sum of 7 on two fair six sided dice B. drawing a spade out of a standard 52 card deck C. drawing a queen out of a standard 52 card deck given it’s a face card. D. drawing a red card out of a standard 52 card deck E. rolling a 4, 5 or 6 on one fair six sided die

OpenStudy (anonymous):

a) \[\frac{1}{6}\] b)\[\frac{1}{4}\] c)\[\frac{1}{3}\] d)\[\frac{1}{2}\] e)\[\frac{1}{2}\]